# Catalog of outbursts of neutron star LMXBs

### v1

@@ 4 Conclusions
We use the information assembled to make a few new conclusions. For instance, we consider arguments about the distance to 4U 1608-52, and identify periods of prolonged quiescence in AX J1754.2-2754, XMM J174457-2850.3, and XMMU J174716.1-281048 (aka IGR J17464-2811). We then look at the distribution of outbursts with time, and flux, to identify a rough empirical flux completeness limit for X-ray binary outbursts since 1996 of $F_{X}=3\times 10^{-10}$ ergs/s/cm2 (though a number of outbursts have been identified below this limit, and some above have been missed).
Finally, we look at the distribution of the frequency of recorded outbursts from these transients. While four transients average roughly one to three outbursts per year, nearly half (40 of 85 systems) have zero or one recorded outburst between 1996 and 2023. It is certainly possible that many of these have shown fainter outbursts below our detection level during this period. However, relatively stringent limits on outbursts from the nearby transient LMXB Cen X-4 for several decades suggest that there are transient neutron star LMXBs with properties consistent with the intriguing Galactic Center X-ray binaries identified by Hailey et al. (2018).
Swift/XRT lightcurves supplied by the UK Swift Science Centre at the University of Leicester (Evans et al., 2007). RXTE/ASM results provided by the ASM/RXTE teams at MIT and at the RXTE SOF and GOF at NASA’s GSFC. RXTE/PCA results from NASA Goddard Space Flight Center, https://asd.gsfc.nasa.gov/Craig.Markwardt//galscan/main.html. INTEGRAL/JEM-X lightcurves from Kuulkers et al. (2007) using the OSA10.1 software, distributed by the ISDC (Courvoisier et al., 2003). MAXI results provided by RIKEN, JAXA and the MAXI team (Matsuoka et al., 2009). This paper made very extensive use of the Astrophysical Data Service (https://ui.adsabs.harvard.edu), ArXiv (http://arxiv.org), and the Astronomer’s Telegram (http://www.astronomerstelegram.org). CH is supported by NSERC Discovery Grant RGPIN-2023-04264. JZ has been supported by a Mitacs Globalink Research Internship. In Appendix A we give the full table (Table A) of data on neutron star LMXB outbursts, except for the frequent outbursters, which are given in Appendix B. Appendix C tabulates information on bursters without recorded outbursts, and Appendix D tabulates the neutron star X-ray binaries that we consider to be persistent systems (not further discussed in this paper).
@@ Appendix A Outburst details
IGR J00291+5934 2008/8 2008/10 6.1e-10 - 1.7 1.3e36 4.2 Lewis et al. (2010)
IGR J00291+5934 2015/8 2015/8 5.6e-10 1e22 1.7 1.2e36 4.2 De Falco et al. (2017)
@@
GS 0836-429 2003/1 2003/5 1.0e-9 2.7e22 1.5 1.0e37 $<$9.2 Aranzana et al. (2016)
GS 0836-429 2003/9 2004/6 1.3e-9 2.7e22 1.5 1.3e37 $<$9.2 Aranzana et al. (2016)
MAXI J0911-655 2016/2 2023/10 1.7e-10 2.7e21 1.27 1.9e36 9.5 Sanna et al. (2017b); Ng et al. (2021a),MAXI
XTE J0929-314 2002/4 2002/4 6.7e-10 7.6e20 1.7-2 $>$2.9e36 $>$6 Galloway et al. (2002); Juett et al. (2003)
@@
MAXI J1421-613 2014/01 2014/1 1.9e-9 4.8e22 2.1 2e36 3 Serino et al. (2015); Nobukawa et al. (2020)
@@
Cen X-4 1979/05 1979/5 9.6e-8 9e20 - 1.9e37 1.3 Kaluzienski et al. (1980); Kuulkers et al. (2009)
MAXI J1621-501 2017/09 2019/01 3.1e-9 6.1e22 2.7 2.6e37 8.4 Gorgone et al. (2019); Chenevez et al. (2018)
MAXI J1647-227 2012/06 2012/06 9.2e-10 4.3e21 1.5 1.1e37 10 Onodera et al. (2014); Kennea et al. (2012)
MXB 1658-298 1978/3 1979/7 1.9e-9 - - 2.3e37 10 Cominsky et al. (1983)
MXB 1658-298 1999/09 1999/09 6.5e-10 - - 7.8e36 10 Heise et al. (1999); Oosterbroek et al. (2001)
MXB 1658-298 2015/09 2015/09 5.5e-10 2.7e21 - 6.6e36 10 Sharma et al. (2018)
IGR J16597-3704 2017/10 2017/10 2.2e-10 1.5e22 1.46 2.2e36 9.1 Sanna et al. (2018)
XTE J1701-407 2008/6 2011/7 7.3e-10 3.4e22 2.0 3.3e36 6.2 Falanga et al. (2009); Linares et al. (2009)
XTE J1701-407 2011/09 2011/09 3.8e-10 3e22 1.4 1.8e36 6.2 Degenaar et al. (2011a)
XTE J1701-407 2016/08 2017/4 3.5e-10 - - 1.6e36 6.2 Swift/XRT
XTE J1701-407 2018/03 2018/03 3.5e-10 - - 1.6e36 6.2 Barthelmy et al. (2018),Swift/XRT
XTE J1701-407 2019/04 2019/04 4e-10 3.2e22 1.9 1.8e36 6.2 Iwakiri et al. (2019)
XTE J1701-407 2020/01 2020/03 1.8e-10 - - 8.2e35 6.2 Onori et al. (2020); Palmer et al. (2020)
XTE J1701-462 2005/12 2007/8 4.3e-8 2.7e22 1.9 4e38 8.8 Lin et al. (2009b); Kennea et al. (2006)
XTE J1701-462 2022/9 2023/3 1.4e-8 3e22 - 1.3e38 8.8 Cocchi et al. (2023),Swift/XRT
IGR J17062-6143 2005/12 2022/9* 1.4e-10 2.3e21 - 8.7e35 7.3 Remillard & Levine (2008b); Bult et al. (2021a)
XTE J1709-267 1990/3 1990/8 4.3e-9 - - 3.7e37 8.5 Yamauchi (2010)
XTE J1709-267 1997/1 1997/4 6.7e-9 - - 5.8e37 8.5 Cocchi et al. (1998)
@@
XTE J1709-267 2023/10 2023/11 1.4e-9 - - 1.2e37 8.5 Negoro et al. (2023),MAXI
2S 1711-339 1975/2 1976/9 3.2e-9 1.5e22 2.2 2.2e37 $<$7.5 Cornelisse et al. (2002b)
@@
2S 1711-339 2016/11 2016/11 1.1e-9 - - 7.7e36 $<$7.5 Negoro et al. (2016b)
1M 1716-315 1971/9 1984/6 1.1e-9 2.1e21 2 6.2e36 6.9 Jonker et al. (2007); Makishima et al. (1981)
IGR J17191-2821 2007/3 2007/5 1.8e-9 3.4e21 1.75 2.7e37 $<$11 Altamirano et al. (2010c); Markwardt et al. (2007)
@@
4U 1730-22 1972/8 1972/8 2.9e-9 - - 1.7e37 6.9 Chen et al. (1997)
4U 1730-22 2021/6 2023/8* 2.4e-9 3.9e21 - 1.4e37 6.9 Chen et al. (2023); Bult et al. (2022a)
KS 1731-260 1988/10 2001/2 5.2e-9 1.1e22 - 3.0e37 7 Wijnands et al. (2001b)
XB 1733-30 1980/7 1999/4 3.3e-9 1.8e22 2.1 1.3e37 5.7 Guainazzi et al. (1999); Baumgardt & Vasiliev (2021)
XB 1733-30 2021/4 2023/7* 2e-9 2.2e22 2 7.7e36 5.7 Homan et al. (2023)
Swift J1734.5-3027 2013/9 2013/9 3.6e-10 9.4e21 2.36 2.2e36 7.2 Bozzo et al. (2015)
IGR J17379-3747 2004/2 2004/2 1.3e-10 - - 1.1e36 8.5 Markwardt et al. (2008); Chelovekov & Grebenev (2010)
@@
XTE J1739-285 2019/9 2020/9 4.5e-9 1.7e22 - 2.9e37 $<$7.3 Negoro (2020); Beri et al. (2023), MAXI
@@
GRS 1741.9-2853 2007/3 2007/4 2.6e-10 - - 2.5e36 9.0 Degenaar & Wijnands (2009, 2010); Trap et al. (2009)
GRS 1741.9-2853 2009/10 2009/10 2.2e-9 - - 2.1e37 9.0 Degenaar & Wijnands (2010)
@@
1A 1744-361 2004/04 2004/04 3.4e-10 - - 2.6e36 8 Markwardt & Swank (2004b)
1A 1744-361 2005/07 2005/07 5.1e-10 - - 3.9e36 8 Swank & Markwardt (2005)
1A 1744-361 2008/06 2008/6 2.2e-09 - - 1.7e37 8 Remillard & Levine (2008a)
@@
1A 1744-361 2010/03 2010/3 5.8e-10 - - 4.4e36 8 RXTE/PCA
1A 1744-361 2011/09 2011/9 1.7e-09 - - 1.3e37 8 RXTE/PCA
1A 1744-361 2013/08 2013/8 2.1e-09 5.1e21 1.5 1.6e37 8 Bahramian et al. (2013b)
1A 1744-361 2022/5 2022/8 2.4e-09 - - 1.9e37 8 Hughes et al. (2022); Ng et al. (2023)
IGR J17445-2747 2004/2 2004/3 7.4e-11 - 2.2 5.7e35 8 Mereminskiy et al. (2017)
IGR J17445-2747 2017/04 2017/4 9.2e-11 4.9e22 2.2 7.1e35 8 Heinke et al. (2017); Mereminskiy et al. (2017)
IGR J17445-2747 2019/05 2019/5 4.6e-12 - 2.2 3.6e34 8 Heinke et al. (2019a)
@@
AX J1745.6-2901 2023/6 2023/11* 3.4e-11 2.2e23 1.0 2.6e35 8 Reynolds et al. (2023)
GRO J1744-28 1995/12 1996/5 3.5e-8 3.5e22 1.2 6.7e37 4 Giles et al. (1996)
GRO J1744-28 1996/12 1997/4 2.1e-8 - - 4.0e37 4 Woods et al. (1999)
GRO J1744-28 2014/1 2014/1 9.7e-9 8.6e22 0.99 1.9e37 4 Degenaar et al. (2014a)
GRO J1744-28 2017/2 2017/2 4.4e-10 7e22 1.2 8.4e35 4 König et al. (2020); Sanna et al. (2017a)
GRO J1744-28 2021/6 2021/ 1.2e-9 9e22 1.1 2.3e36 4 Mereminskiy et al. (2021)
EXO 1745-248 2000/7 2000/8 1.4e-8 - - 6e37 5.9 Heinke et al. (2003a)
EXO 1745-248 2011/10 2011/11 2.1e-9 1.2e22 2.25 8.8e36 5.9 Serino et al. (2012)
EXO 1745-248 2015/3 2015/6 2.0e-8 3.0e22 0.93 8.4e37 5.9 Tetarenko et al. (2016a)
IGR J17480-2446 2010/10 2010/11 1.4e-8 1.2e22 2.1 5.1e37 5.9 Patruno et al. (2012); Miller et al. (2011)
Swift J174805.3-244637 2012/7 2012/8 1.1e-8 - 1.6 4.6e37 5.9 Bahramian et al. (2014)
Swift J174805.3-244637 2023/2 2023/3 2.5e-9 $2$ 1.0e37 5.9 Heinke et al. (2023); Sanna et al. (2023a)
IGR J17473-2721 2005/5 2005/5 1.2e-9 - - 4.3e36 5.5 Markwardt & Swank (2005a)
IGR J17473-2721 2008/3 2008/8 1.3e-8 3.8e22 - 4.7e37 5.5 Altamirano et al. (2008b); Chenevez et al. (2011b)
XMMU J174716.1-281048 2003/3 2012/10 7.4e-12 9e22 2.25 6.3e34 $<$8.4 Degenaar & Wijnands (2007); Kaur et al. (2017)
@@
SAX J1747.0-2853 2017/4 2017/4 5.4e-10 - - 4e36 7.5 Heinke et al. (2017)
SAX J1747.0-2853 2019/2 2019/2 1.3e-8 - - 9e37 7.5 Negoro et al. (2019); Mereminskiy et al. (2019)
SAX J1747.0-2853 2022/4 2022/5 7e-9 - - 5e37 7.5 Iwakiri et al. (2022); Kennea et al. (2022b)
SAX J1747.0-2853 2023/5 2023/6* 1.4e-9 - - 9.4e36 7.5 Li et al. (2023),MAXI,Swift/XRT
SAX J1748.9-2021? 1971/12 1972/1 7.3e-9 5.3e22 2.1 6.3e37 8.5 Markert et al. (1975); Forman et al. (1976)
SAX J1748.9-2021 1998/8 1998/8 1.3e-9 8.2e21 1.44 1.1e37 8.5 in ’t Zand et al. (1999)
@@
SAX J1748.9-2021 2015/2 2015/4 4e-9 5.7e21 2.3 3.4e37 8.5 Pintore et al. (2016)
SAX J1748.9-2021 2017/10 2017/10 1.9e-10 5.1e21 1.6 1.6e36 8.5 Pintore et al. (2018)
SAX J1748.9-2021 2021/11 2021/12 4.9e-10 6e22 1.8 4.2e36 8.5 Pike et al. (2021a, b)
IGR J17494-3030 2012/3 2012/3 1.0e-10 1.8e22 1.9 7.7e35 8 Armas Padilla et al. (2013)
@@
Swift J1749.4-2807 2006/6 2006/6 9.5e-11 3e22 5e35 6.7 Wijnands et al. (2009)
Swift J1749.4-2807 2010/4 2010/4 5.2e-10 - - 2.8e36 6.7 Ferrigno et al. (2011); Altamirano et al. (2011b)
Swift J1749.4-2807 2021/3 2021/3 3.1e-10 4.5e22 - 1.7e36 6.7 Marino et al. (2022)
IGR J17498-2921 2011/8 2011/9 5.1e-10 1.2e22 1.8 3.5e36 7.6 Falanga et al. (2012); Linares et al. (2011)
IGR J17498-2921 2023/4 2023/5 6.1e-10 1.2e22 1.8 4.2e36 7.6 Sanna et al. (2023b), NICER
SAX J1750.8-2900 1997/3 1997/3 3.5e-9 6e22 2.4 1.9e37 6.79 Natalucci et al. (1999)
@@
SAX J1750.8-2900 2008/3 2008/10 5.1e-9 2.3e22 1.2 2.8e37 6.79 Markwardt & Swank (2008); Linares et al. (2008)
SAX J1750.8-2900 2011/2 2011/2 3.6e-10 3.2e22 2 2.0e36 6.79 Natalucci et al. (2011)
SAX J1750.8-2900 2015/9 2016/3 2.4e-9 2.3e22 2 1.3e37 6.79 Sanchez-Fernandez et al. (2015), IGR/JEM-X
@@
XTE J1751-305 1998/6 1998/6 1.7e-9 - 1.8 1.5e37 8.5 Markwardt et al. (2002)
XTE J1751-305 2002/4 2002/4 1.1e-9 - 1.8 9.9e36 8.5 Markwardt et al. (2002)
XTE J1751-305 2005/3 2005/3 1.8e-10 - - 1.6e36 8.5 Swank et al. (2005)
@@
XTE J1751-305 2009/10 2009/10 3.2e-10 9.8e21 - 2.8e36 8.5 Chenevez et al. (2009); Miller et al. (2003b)
IGR J17511-3057 2009/9 2009/9 4.5e-10 5e21 1.2 2.6e36 6.9 Bozzo et al. (2010)
IGR J17511-3057 2015/3 2015/4 5e-10 1.2e22 1.5 2.8e36 6.9 Papitto et al. (2016)

### v2

@@ 4 Conclusions
We use the information assembled to make a few new conclusions. For instance, we consider arguments about the distance to 4U 1608-52, and identify periods of prolonged quiescence in AX J1754.2-2754, XMM J174457-2850.3, and XMMU J174716.1-281048 (aka IGR J17464-2811). We then look at the distribution of outbursts with time, and flux, to identify a rough empirical flux completeness limit for X-ray binary outbursts since 1996 of $F_{X}=3\times 10^{-10}$ ergs cm-2 s-1 (though a number of outbursts have been identified below this limit, and some above have been missed).
Finally, we look at the distribution of the frequency of recorded outbursts from these transients. While four transients average roughly one to three outbursts per year, nearly half (40 of 87 systems) have zero or one recorded outburst between 1996 and 2023. It is certainly possible that many of these have shown fainter outbursts below our detection level during this period. However, relatively stringent limits on outbursts from nearby transient LMXBs, such as Cen X-4, PSR J1023+0038, and PSR J12270-4859, for several decades suggest that there are transient neutron star LMXBs with properties consistent with the intriguing Galactic Center X-ray binaries identified by Hailey et al. (2018).
Note added in proof: Ni et al. (2025) report a new X-ray transient detected by the Einstein Probe. Wu et al. (2025a) report a refined position, an eclipse, and identify two flares in its X-ray lightcurve as X-ray bursts, indicating that EP J171159.4-333253 is almost certainly a new neutron star X-ray transient.
We thank the referee for an exceptionally thoughtful and useful report. Swift/XRT lightcurves supplied by the UK Swift Science Centre at the University of Leicester (Evans et al., 2007). RXTE/ASM results provided by the ASM/RXTE teams at MIT and at the RXTE SOF and GOF at NASA’s GSFC. RXTE/PCA results from NASA Goddard Space Flight Center, https://asd.gsfc.nasa.gov/Craig.Markwardt//galscan/main.html. INTEGRAL/JEM-X lightcurves from Kuulkers et al. (2007a) using the OSA10.1 software, distributed by the ISDC (Courvoisier et al., 2003). MAXI results provided by RIKEN, JAXA and the MAXI team (Matsuoka et al., 2009). This paper made very extensive use of the Astrophysical Data Service (https://ui.adsabs.harvard.edu), ArXiv (http://arxiv.org), and the Astronomer’s Telegram (http://www.astronomerstelegram.org). CH is supported by NSERC Discovery Grant RGPIN-2023-04264. JZ has been supported by a Mitacs Globalink Research Internship. This paper employs a list of Chandra datasets, obtained by the Chandra X-ray Observatory, contained in the Chandra Data Collection (CDC) 346: https://doi.org/10.25574/cdc.346 (catalog doi:10.25574/cdc.346). In Appendix A we give the full table (Table A) of data on neutron star LMXB outbursts, except for the frequent outbursters, which are given in Appendix B. Appendix C tabulates information on bursters without recorded outbursts, and Appendix D tabulates the neutron star X-ray binaries that we consider to be persistent systems (not further discussed in this paper).
@@ Appendix A Outburst details
IGR J00291+5934 2008/8 2008/10 6.1e-10 - 1.7 1.3e36 4.2 Chakrabarty et al. (2008); Lewis et al. (2010)
IGR J00291+5934 2015/8 2015/8 5.6e-10 1e22 1.7 1.2e36 4.2 Cummings et al. (2015a); De Falco et al. (2017)
@@
GS 0836-429 2003/1 2003/5 1.0e-9 2.7e22 1.5 1.0e37 $<$9.2 Rodriguez et al. (2003); Aranzana et al. (2016)
GS 0836-429 2003/9 2004/6 1.3e-9 2.7e22 1.5 1.3e37 $<$9.2 Chelovekov et al. (2005); Aranzana et al. (2016)
MAXI J0911-655 2016/2 2023/10 1.7e-10 2.7e21 1.27 1.9e36 9.5 Serino et al. (2016b); Sanna et al. (2017b); Ng et al. (2021a)
XTE J0929-314 2002/4 2002/4 6.7e-10 7.6e20 1.7-2 $>$2.9e36 $>$6 Remillard (2002); Galloway et al. (2002); Juett et al. (2003)
@@
MAXI J1421-613 2014/1 2014/1 1.9e-9 4.8e22 2.1 2e36 3 Morooka et al. (2014); Serino et al. (2015); Nobukawa et al. (2020)
@@
Cen X-4 1979/5 1979/5 9.6e-8 9e20 - 1.9e37 1.3 Kaluzienski et al. (1980); Kuulkers et al. (2009)
MAXI J1621-501 2017/9 2019/1 3.1e-9 6.1e22 2.7 2.6e37 8.4 Hashimoto et al. (2017); Gorgone et al. (2019); Chenevez et al. (2018)
MAXI J1647-227 2012/6 2012/6 9.2e-10 4.3e21 1.5 1.1e37 10 Negoro et al. (2012); Onodera et al. (2014); Kennea et al. (2012)
MXB 1659-298 1976/10 1979/7 1.9e-9 - - 2.3e37 10 Lewin et al. (1976a); Cominsky et al. (1983)
MXB 1659-298 1999/9 1999/9 6.5e-10 - - 7.8e36 10 in ’t Zand et al. (1999a); Heise et al. (1999); Oosterbroek et al. (2001)
MXB 1659-298 2015/9 2015/9 5.5e-10 2.7e21 - 6.6e36 10 Negoro et al. (2015a); Sharma et al. (2018)
IGR J16597-3704 2017/10 2017/10 2.2e-10 1.5e22 1.46 2.2e36 9.1 Bozzo et al. (2017); Sanna et al. (2018)
XTE J1701-407 2008/6 2011/7 7.3e-10 3.4e22 2.0 3.3e36 6.2 Markwardt et al. (2008c); Falanga et al. (2009); Linares et al. (2009a)
XTE J1701-407 2011/9 2011/9 3.8e-10 3e22 1.4 1.8e36 6.2 Degenaar et al. (2011a)
XTE J1701-407 2016/8 2017/4 3.5e-10 - - 1.6e36 6.2 Swift/XRT
XTE J1701-407 2018/3 2018/3 3.5e-10 - - 1.6e36 6.2 Barthelmy et al. (2018),Swift/XRT
XTE J1701-407 2019/4 2019/4 4e-10 3.2e22 1.9 1.8e36 6.2 Iwakiri et al. (2019)
XTE J1701-407 2020/1 2020/3 1.8e-10 - - 8.2e35 6.2 Onori et al. (2020); Palmer et al. (2020)
XTE J1701-462 2005/12 2007/8 4.3e-8 2.7e22 1.9 4e38 8.8 Remillard et al. (2006); Lin et al. (2009b); Kennea et al. (2006)
XTE J1701-462 2022/9 2023/3 1.4e-8 3e22 - 1.3e38 8.8 Iwakiri et al. (2022a); Cocchi et al. (2023),Swift/XRT
IGR J17062-6143 2005/12 2022/9* 1.4e-10 2.3e21 - 8.7e35 7.3 Churazov et al. (2007); Remillard & Levine (2008b); Bult et al. (2021a)
XTE J1709-267 1990/3 1990/8 4.3e-9 - - 3.7e37 8.5 Yamauchi & Koyama (1990); Yamauchi (2010)
XTE J1709-267 1997/1 1997/4 6.7e-9 - - 5.8e37 8.5 Marshall et al. (1997); Cocchi et al. (1998)
@@
XTE J1709-267 2023/10 2023/11 1.4e-9 - - 1.2e37 8.5 Negoro et al. (2023a),MAXI
XTE J1709-267 2025/5 2025/6* 3.5e-10 3.9e21 2.0 3.0e36 8.5 Wu et al. (2025b)
2S 1711-339 1975/2 1976/9 3.2e-9 1.5e22 2.2 2.2e37 $<$7.5 Carpenter et al. (1977); Cornelisse et al. (2002b)
@@
2S 1711-339 2016/11 2016/11 1.1e-9 - - 7.7e36 $<$7.5 Negoro et al. (2016c)
1M 1716-315 1971/9 1984/6 1.1e-9 2.1e21 2 6.2e36 6.9 Markert et al. (1976); Makishima et al. (1981a); Jonker et al. (2007)
IGR J17191-2821 2007/3 2007/5 1.8e-9 3.4e21 1.75 2.7e37 $<$11 Turler et al. (2007); Altamirano et al. (2010c); Markwardt et al. (2007)
@@
4U 1730-22 1972/8 1972/8 2.9e-9 - - 1.7e37 6.9 Cominsky et al. (1978); Chen et al. (1997)
4U 1730-22 2021/6 2024/3 2.4e-9 3.9e21 - 1.4e37 6.9 Kobayashi et al. (2021); Chen et al. (2023a); Bult et al. (2022a), MAXI, Swift/XRT
KS 1731-260 1988/10 2001/2 5.2e-9 1.1e22 - 3.0e37 7 Syunyaev et al. (1990); Wijnands et al. (2001b)
XB 1733-30 1980/7 1999/4 3.3e-9 1.8e22 2.1 1.3e37 5.7 Makishima et al. (1981b); Guainazzi et al. (1999)
XB 1733-30 2020/4 2025/5* 2e-9 2.2e22 2 7.7e36 5.7 Mereminskiy et al. (2020b); Negoro et al. (2023b); Homan et al. (2023)
Swift J1734.5-3027 2013/5 2013/6 3e-10 - - 2e36 7.2 La Parola et al. (2013); Bozzo et al. (2015a)
Swift J1734.5-3027 2013/9 2013/9 3e-10 1e22 1.7 1.9e36 7.2 Kennea et al. (2013); Bozzo et al. (2015a)
Swift J1734.5-3027 2013/10 2014/4 3.6e-10 1e22 2.36 2.2e36 7.2 Bozzo et al. (2015a)
IGR J17379-3747 2004/2 2004/2 1.3e-10 - - 1.1e36 8.5 Markwardt et al. (2008a); Chelovekov & Grebenev (2010)
@@
XTE J1739-285 2019/9 2020/9 4.5e-9 1.7e22 - 2.9e37 $<$7.3 Mereminskiy & Grebenev (2019); Negoro (2020); Beri et al. (2023), MAXI
@@
GRS 1741.9-2853 2007/3 2007/4 2.6e-10 - - 2.5e36 9.0 Kuulkers et al. (2007b); Degenaar & Wijnands (2009); Trap et al. (2009)
GRS 1741.9-2853 2009/10 2009/10 2.2e-9 - - 2.1e37 9.0 Chenevez et al. (2009a); Degenaar & Wijnands (2010)
@@
1A 1744-361 2004/4 2004/4 3.4e-10 - - 2.6e36 8 Markwardt & Swank (2004b)
1A 1744-361 2005/7 2005/7 5.1e-10 - - 3.9e36 8 Swank & Markwardt (2005)
1A 1744-361 2008/6 2008/6 2.2e-09 - - 1.7e37 8 Remillard & Levine (2008a)
@@
1A 1744-361 2010/3 2010/3 5.8e-10 - - 4.4e36 8 RXTE/PCA
1A 1744-361 2011/9 2011/9 1.7e-09 - - 1.3e37 8 RXTE/PCA
1A 1744-361 2013/8 2013/8 2.1e-09 5.1e21 1.5 1.6e37 8 Bahramian et al. (2013b)
1A 1744-361 2022/5 2022/8 2.4e-09 - - 1.9e37 8 Kobayashi et al. (2022); Hughes et al. (2022); Ng et al. (2023)
1A 1744-361 2024/7 2025/2* 5.9e-9 5.7e21 1.4 4.5e37 8 Xu et al. (2024); Cheng et al. (2024)
IGR J17445-2747 2004/2 2004/3 7.4e-11 - 2.2 5.7e35 8 Bird et al. (2006); Mereminskiy et al. (2017b)
IGR J17445-2747 2017/4 2017/4 9.2e-11 4.9e22 2.2 7.1e35 8 Heinke et al. (2017); Mereminskiy et al. (2017b)
IGR J17445-2747 2019/5 2019/5 4.6e-12 - 2.2 3.6e34 8 Heinke et al. (2019a)
@@
AX J1745.6-2901 2023/6 2025/5* 3.4e-11 2.2e23 1.0 2.6e35 8 Reynolds et al. (2023); Degenaar et al. (2025)
GRO J1744-28 1995/12 1996/5 3.5e-8 3.5e22 1.2 6.7e37 4 Fishman et al. (1995); Giles et al. (1996)
GRO J1744-28 1996/12 1997/4 2.1e-8 - - 4.0e37 4 Kouveliotou et al. (1997); Woods et al. (1999)
GRO J1744-28 2014/1 2014/1 9.7e-9 8.6e22 0.99 1.9e37 4 Negoro et al. (2014a); Degenaar et al. (2014a)
GRO J1744-28 2017/2 2017/2 4.4e-10 7e22 1.2 8.4e35 4 Mereminskiy et al. (2017a); König et al. (2020); Sanna et al. (2017a)
GRO J1744-28 2021/6 2021/6 1.2e-9 9e22 1.1 2.3e36 4 Mereminskiy et al. (2021a)
EXO 1745-248 2000/7 2000/8 1.4e-8 - - 6e37 5.9 Markwardt & Swank (2000); Heinke et al. (2003a)
EXO 1745-248 2011/10 2011/11 2.1e-9 1.2e22 2.25 8.8e36 5.9 Altamirano et al. (2011a); Serino et al. (2012)
EXO 1745-248 2015/3 2015/6 2.0e-8 3.0e22 0.93 8.4e37 5.9 Altamirano et al. (2015); Tetarenko et al. (2016a)
IGR J17480-2446 2010/10 2010/11 1.4e-8 1.2e22 2.1 5.1e37 5.9 Bordas et al. (2010); Patruno et al. (2012); Miller et al. (2011)
Swift J174805.3-244637 2012/7 2012/8 1.1e-8 - 1.6 4.6e37 5.9 Wijnands et al. (2012); Bahramian et al. (2014)
Swift J174805.3-244637 2023/2 2023/3 2.5e-9 $2$ 1.0e37 5.9 Negoro et al. (2023c); Heinke et al. (2023); Sanna et al. (2023a)
IGR J17473-2721 2005/5 2005/5 1.2e-9 - - 4.3e36 5.5 Grebenev et al. (2005b); Markwardt & Swank (2005a)
IGR J17473-2721 2008/3 2008/8 1.3e-8 3.8e22 - 4.7e37 5.5 Del Monte et al. (2008); Altamirano et al. (2008b); Chenevez et al. (2011b)
XMMU J174716.1-281048 2003/3 2012/10 7.4e-12 9e22 2.25 6.3e34 $<$8.4 Sidoli & Mereghetti (2003); Degenaar & Wijnands (2007); Kaur et al. (2017)
@@
SAX J1747.0-2853 2017/4 2017/10 5.4e-10 - - 4e36 7.5 Heinke et al. (2017),Swift/XRT
SAX J1747.0-2853 2019/2 2019/9 1.3e-8 - - 9e37 7.5 Negoro et al. (2019); Mereminskiy et al. (2019),Swift/XRT
SAX J1747.0-2853 2022/4 2022/8 7e-9 - - 5e37 7.5 Iwakiri et al. (2022b); Kennea et al. (2022b),Swift/XRT
SAX J1747.0-2853 2023/5 2023/6 1.4e-9 - - 9.4e36 7.5 Negoro et al. (2023d); Li et al. (2023),MAXI,Swift/XRT
SAX J1748.9-2021? 1971/12 1972/1 7.3e-9 5.3e22 2.1 6.3e37 8.5 Markert et al. (1975); Forman et al. (1976a)
SAX J1748.9-2021 1998/8 1998/8 1.3e-9 8.2e21 1.44 1.1e37 8.5 in ’t Zand et al. (1998b, 1999b)
@@
SAX J1748.9-2021 2015/2 2015/4 4e-9 5.7e21 2.3 3.4e37 8.5 Kuulkers et al. (2015); Pintore et al. (2016)
SAX J1748.9-2021 2017/10 2017/10 1.9e-10 5.1e21 1.6 1.6e36 8.5 Negoro et al. (2017b); Pintore et al. (2018)
SAX J1748.9-2021 2021/11 2021/12 4.9e-10 6e22 1.8 4.2e36 8.5 Negoro et al. (2021b); Pike et al. (2021a, b)
IGR J17494-3030 2012/3 2012/3 1.0e-10 1.8e22 1.9 7.7e35 8 Boissay et al. (2012); Armas Padilla et al. (2013)
@@
Swift J1749.4-2807 2006/6 2006/6 9.5e-11 3e22 5e35 6.7 Schady et al. (2006); Wijnands et al. (2009)
Swift J1749.4-2807 2010/4 2010/4 5.2e-10 - - 2.8e36 6.7 Pavan et al. (2010); Ferrigno et al. (2011); Altamirano et al. (2011c)
Swift J1749.4-2807 2021/3 2021/3 3.1e-10 4.5e22 - 1.7e36 6.7 Mereminskiy et al. (2021b); Marino et al. (2022)
IGR J17498-2921 2011/8 2011/9 5.1e-10 1.2e22 1.8 2.0e36 5.7 Gibaud et al. (2011); Falanga et al. (2012); Linares et al. (2011b)
IGR J17498-2921 2023/4 2023/5 6.1e-10 1.2e22 1.8 2.4e36 5.7 Grebenev et al. (2023); Sanna et al. (2023b), NICER
SAX J1750.8-2900 1997/3 1997/3 3.5e-9 6e22 2.4 1.9e37 6.79 Bazzano et al. (1997); Natalucci et al. (1999)
@@
SAX J1750.8-2900 2008/3 2008/10 5.1e-9 2.3e22 1.2 2.8e37 6.79 Markwardt & Swank (2008a); Linares et al. (2008)
SAX J1750.8-2900 2011/2 2011/2 3.6e-10 3.2e22 2 2.0e36 6.79 Fiocchi et al. (2011); Natalucci et al. (2011)
SAX J1750.8-2900 2015/9 2016/3 2.4e-9 2.3e22 2 1.3e37 6.79 Cummings et al. (2015b); Sanchez-Fernandez et al. (2015), IGR/JEM-X
@@
XTE J1751-305 1998/6 1998/6 1.7e-9 - 1.8 1.5e37 8.5 Markwardt et al. (2002b)
XTE J1751-305 2002/4 2002/4 1.1e-9 - 1.8 9.9e36 8.5 Markwardt & Swank (2002); Markwardt et al. (2002b)
XTE J1751-305 2005/3 2005/3 1.8e-10 - - 1.6e36 8.5 Grebenev et al. (2005c); Swank et al. (2005)
@@
XTE J1751-305 2009/10 2009/10 3.2e-10 9.8e21 - 2.8e36 8.5 Chenevez et al. (2009a); Miller et al. (2003b)
IGR J17511-3057 2009/9 2009/9 4.5e-10 5e21 1.2 2.6e36 6.9 Baldovin et al. (2009); Bozzo et al. (2010)
IGR J17511-3057 2015/3 2015/4 5e-10 1.2e22 1.5 2.8e36 6.9 Bozzo et al. (2015b); Papitto et al. (2016)
IGR J17511-3057 2025/2 2025/2* 3e-10 - - 2e36 6.9 Sguera & Sidoli (2025); Sguera (2025)